# Twisted compactifications and conformal defects from ISO(3)$\times$×U(1) $F(4)$ gauged supergravity

### v1

@@ Abstract
We study holographic solutions describing RG flows across dimensions from five-dimensional $N=2$ SCFT to SCFTs in three and two dimensions using matter-coupled $F(4)$ gauged supergravity with $ISO(3)\times U(1)$ gauge group. By performing topological twists implemented by $SO(2)$, $SO(2)\times SO(2)$ and $SO(3)$ gauge fields, we find a number of $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ preserving eight and four supercharges. These geometries are identified with conformal fixed points in three and two dimensions, respectively. The corresponding flow solutions are obtained numerically and can be interpreted as black membranes and black strings in asymptotically $AdS_{6}$ space. In addition, we have constructed charged domain wall solutions with $AdS_{2}\times S^{3}$ and $AdS_{3}\times S^{2}$ slicing. These solutions are supported by a non-vanishing two-form field and describe conformal line and surface defects within the $N=2$ SCFT in five dimensions, respectively. All of these solutions can be uplifted to type IIB theory via consistent truncations on $S^{2}\times\Sigma$ with $\Sigma$ being a Riemann surface.
@@ 1 Introduction
Five-dimensional field theories possess interesting non-trivial conformal fixed points at infinite coupling constants [1, 2, 3]. These fixed points are interesting in their own right due to the fact that they realize the only possible $N=2$ superconformal symmetry in five dimensions [4]. The essentially strongly-coupled nature and the lack of a Lagrangian formulation make understanding these field theories and the corresponding fixed points a non-trivial task. Up to now, it is well-known that the AdS/CFT correspondence [5, 6, 7] offers a new duality to gain more insights into the dynamics of strongly-coupled field theories via gravity duals. Studies along this line have led to many interesting results, see for example [8]-[43]. These results, giving rise to various insights into different aspects of the five-dimensional superconformal field theories (SCFTs), have been obtained from both ten-dimensional string theories and effective six-dimensional gauged supergravities. Among these, holographic descriptions of twisted compactifications to lower-dimensional SCFTs [30, 29, 33, 34, 35, 36], as initiated in [44], and different types of conformal defects [37, 38, 39, 40, 41, 42, 43] have recently attracted much attention.
In this paper, we are interested in solutions describing twisted compactifications of five-dimensional $N=2$ SCFT to three and two dimensions and conformal line and surface defects in the framework of six-dimensional $N=(1,1)$ gauged supergravity, usually called $F(4)$ gauged supergravity. All the previous works on these types of solutions have considered only pure $F(4)$ gauged supergravity [45] with $SU(2)$ gauge group or a simple extension to matter-coupled theory with $SU(2)\times SU(2)$ gauge group [46, 47]. The former can be embedded in massive type IIA theory using the result of [48] while the higher-dimensional origin of the latter is currently unknown. This renders many interesting hologrpahic solutions including a non-trivial $AdS_{6}$ vacuum with $SU(2)_{\textrm{diag}}$ symmetry identified in [27] not fully useful in the holographic context due to the lack of embedding in string/M-theory. We will partially fill this gap by considering $F(4)$ gauged supergravity coupled to four vector multiplets with $ISO(3)\times U(1)$ gauge group. This gauged supergravity and a truncation to three vector multiplets with $ISO(3)$ gauge group have been shown in [49] to arise from consistent truncations of type IIB theory on $S^{2}\times\Sigma$ for a Riemann surface $\Sigma$.
@@
For solutions describing conformal defects within $N=2$ SCFT in five dimensions, we will study solutions dual to line and surface defects preserving conformal symmetry in one and two dimensions. These solutions will be obtained by taking the metric ansatz in the form of $AdS_{2}\times S^{3}$- and $AdS_{3}\times S^{2}$-sliced domain walls. As we will see, the results found here and in [50] indicate that many aspects of the solutions are very similar to those previously found in compact $SU(2)$ and $SU(2)\times SU(2)$ gauge groups. Although the structure of solutions is not rich as in the case of compact $SU(2)\times SU(2)$ gauge group due to the lack of a non-trivial $AdS_{6}$ vacuum in addition to the trivial one at the origin of the scalar manifold, many properties and even the form of the solutions are very similar. This is in agreement with the general expectation that solutions of gauged supergravity that has no known higher-dimensional origins might still capture some insights into the general aspect of strongly-coupled dynamics of the dual field theories.
The paper is organized as follows. In section 2, we review the general structure of matter-coupled $F(4)$ gauged supergravity as constructed in [46, 47]. We then consider the case of four vector multiplets and $ISO(3)\times U(1)$ gauge group and present the supersymmetric $AdS_{6}$ vacuum studied in [50]. In section 3, we study supersymmetric $AdS_{4}\times\Sigma$ and $AdS_{3}\times\mathcal{M}_{3}$ solutions using the domain wall ansatz with $\mathbb{R}^{1,2}\times\Sigma$ and $\mathbb{R}^{1,1}\times\mathcal{M}_{3}$ slicing and implementing topological twists by turning on $SO(2)\times U(1)$ and $SO(3)$ gauge fields. A number of $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ including numerical solutions interpolating between these geometries and the supersymmetric $AdS_{6}$ vacuum will also be given. We then move to the study of solutions describing conformal line and surface defects in section 4. In this case, the metric ansatz takes the forms of $AdS_{2}\times S^{3}$- and $AdS_{3}\times S^{2}$-sliced domain walls, respectively. Unlike the case of twisted compactifications in which the two-form field from the supergravity multiplet can be consistently truncated out, we consider solutions with a non-vanishing two-form field and set all the gauge fields to zero. We find solutions describing line and surface defects due to a source of a dimension-$4$ operator dual to the two-form field. Finally, we give some conclusions and comments in section 5.
@@ 2 Matter-coupled $F(4)$ gauged supergravity with $ISO(3)\times U(1)$ gauge group
$e^{-1}\mathcal{L}$ $=$ $\frac{1}{4}R-e\partial_{\mu}\sigma\partial^{\mu}\sigma-\frac{1}{4}P^{I\alpha}_{\mu}P^{\mu}_{I\alpha}-\frac{1}{8}e^{-2\sigma}\mathcal{N}_{\Lambda\Sigma}\widehat{F}^{\Lambda}_{\mu\nu}\widehat{F}^{\Sigma\mu\nu}-\frac{3}{64}e^{4\sigma}H_{\mu\nu\rho}H^{\mu\nu\rho}$
@@
$\widehat{F}^{\Lambda}=F^{\Lambda}-m\delta^{\Lambda 0}B,\qquad F^{\Lambda}=dA^{\Lambda}-\frac{1}{2}f^{\phantom{\Sigma}\Lambda}_{\Sigma\phantom{\Lambda}\Gamma}A^{\Sigma}\wedge A^{\Gamma},\qquad H=dB\,.$ (5)
$f^{\phantom{\Sigma}\Lambda}_{\Sigma\phantom{\Lambda}\Gamma}$ are structure constants of the gauge group. It is also useful to note the convention on differential forms used in [46, 47]
@@
and $H_{\mu\nu\rho}=\frac{1}{3}(\partial_{\mu}B_{\nu\rho}+\partial_{\nu}B_{\rho\mu}+\partial_{\rho}B_{\mu\nu})$ (7)
@@
with ${\sigma^{rA}}_{B}$ being Pauli matrices. $SU(2)_{R}$ fundamental indices $A,B$ are raised and lowered by $\epsilon^{AB}$ and $\epsilon_{AB}$ with the convention $T^{A}=\epsilon^{AB}T_{B}$ and $T_{A}=T^{B}\epsilon_{BA}$. The covariant derivative of $\epsilon_{A}$ is defined by
@@
$Q_{AB}=\frac{i}{2}\sigma^{r}_{AB}\left[\frac{1}{2}\epsilon^{rst}\Omega_{\mu st}-i\gamma_{7}\Omega_{\mu r0}\right].$ (19)
In this paper, we are interested in the case of $n=4$ vector multiplets leading to +×$\mathbb{R}^{+}\times SO(4,4)$ global symmetry. The gauge group of interest here is $ISO(3)\times U(1)$ with $ISO(3)$ embedded in $SO(4,3)\subset SO(4,4)$ and the $U(1)$ factor is the abelian gauge symmetry associated with the fourth vector multiplet as in the ungauged theory. The structure constants corresponding to the $ISO(3)\sim SO(3)\ltimes\mathbf{R}^{3}$ factor are given by
@@
with indices $I,J=1,2,3,4$ split as $I=(\bar{r},4)$ for $\bar{r},\bar{s}=1,2,3$. The resulting gauge generators are given by $X_{\Lambda}=(X_{0},X_{r},X_{\bar{r}})$ with $X_{0}=0$, and $X_{r}$ and $X_{\bar{r}}$ respectively generate the $SO(3)$ compact subgroup and three abelian translations $\mathbf{R}^{3}$ transforming as adjoint representation of $SO(3)$, see more detail in [50].
This gauged supergravity admits a supersymmetric $AdS_{6}$ vacuum at the origin of the $SO(4,4)/SO(4)\times SO(4)$ scalar manifold. The $AdS_{6}$ vacuum is given by, for more detail see [50],
@@
$V_{0}$ and $L$ denote the cosmological constant and the $AdS_{6}$ radius, respectively. As usual, we can choose $g=3m$ by shifting the value of the dilaton to $\sigma=0$ at the vacuum. This results in
@@
in which we have chosen $m>0$.
@@ 3 Twisted compactifications of $N=2$ SCFT from ISO(3)$\times$×U(1) $F(4)$ gauged supergravity
In this section, we consider supersymmetric $AdS_{4}\times\Sigma$ and $AdS_{3}\times\mathcal{M}_{3}$ with $\Sigma$ and $\mathcal{M}_{3}$ being a Riemann surface and a constant curvature $3$-manifold, respectively. These solutions are dual to conformal fixed points in three and two dimensions of the five-dimensional $N=2$ SCFT dual to the supersymmetric $AdS_{6}$ vacuum mentioned in the previous section. The solutions interpolating between these geometries and the $AdS_{6}$ vacuum are then interpreted as RG flows across dimensions from five dimensions to these fixed points in lower dimensions. In order to preserve some amount of supersymmetry, we need to perform a topological twist along $\Sigma$ and $\mathcal{M}_{3}$ by turning on suitable gauge fields to cancel spin connections on these manifolds. The lower-dimensional fixed points then arise from twisted compactifications of the $N=2$ SCFT in five dimensions on $\Sigma$ and $\mathcal{M}_{3}$. Moreover, the solutions can also be interpreted as black membranes and black strings in asymptotically $AdS_{6}$ space with near horizon geometries given by $AdS_{4}\times\Sigma$ and $AdS_{3}\times\mathcal{M}_{3}$.
@@ 3.1 Compactifications to three dimensions
$ds^{2}=e^{2f(r)}dx_{1,2}^{2}+dr^{2}+e^{2h(r)}(d\theta^{2}+F_{\kappa}(\theta)d\phi^{2})$ (23)
with $dx^{2}_{1,2}$ being a flat Minkowski metric in three dimensions. The function $F_{\kappa}(\theta)$ is defined by
@@
By denoting the six-dimensional coordinates by $x^{\mu}=(x^{a},r,\theta,\phi)$, for $a=0,1,2$, we can choose the following choice of vielbein
@@
With this ansatz for the gauge fields, it is consistent to set the two-form field to zero. A straightforward computation of the composite connection gives
$Q_{AB}=-\frac{i}{2}gA^{3}\sigma^{3}_{AB}\,.$ (31)
@@
$0=\frac{F^{\prime}_{\kappa}(\theta)}{F_{\kappa}(\theta)}e^{-h}\gamma_{\hat{\phi}\hat{\theta}}\epsilon_{A}-\frac{i}{2}ga\frac{F^{\prime}_{\kappa}(\theta)}{F_{\kappa}(\theta)}e^{-h}\sigma^{3}_{AB}\epsilon^{B}+\ldots\,.$ (32)
We then perform the twist by imposing a projector
@@
We are now ready to analyze the BPS conditions arising from setting supersymmetry transformations of fermions to zero. With only the $(\theta\phi)$-component of the field strength tensors non-vanishing, the field equations of gauge fields require $\phi_{3}=0$. Furthermore, consistency of the BPS equations also require $\phi_{0}=\phi_{4}=0$. By imposing another projector of the form
@@
It can be readily verified that these equations are compatible with the corresponding field equations. The sign choice in (36) has been chosen such that the $AdS_{6}$ vacuum appears at $r\rightarrow\infty$.
@@
It should be pointed out that the fixed points with $\phi_{5}=0$ are possible only for $b=0$ which implies $A^{7}=0$. It is clearly seen that only $\kappa=-1$ is possible leading to an $AdS_{4}\times H^{2}$ solution. This solution can also be interpreted as an $AdS_{4}\times H^{2}$ solution of $F(4)$ gauged supergravity coupled to three vector multiplets with $ISO(3)$ gauge group.
Another $AdS_{2}\times\Sigma$ solution is given by
@@
As in the previous solution, this fixed point exists only for $\kappa=-1$ leading to another $AdS_{4}\times H^{2}$ solution. These solutions should be dual to conformal fixed points in three dimensions arising from twisted compactifications of $N=2$ SCFT in five dimensions on a two-dimensional hyperbolic space $H^{2}$. Since these fixed point solutions are half-BPS, due to the projector (33), the solutions preserve eight supercharges corresponding to $N=2$ superconformal symmetry in three dimensions.
@@
Figure 1: Supersymmetric RG flows from $N=2$ SCFT in five dimensions to $N=2$ SCFT in three dimensions from $ISO(3)$ $F(4)$ gauged supergravity.
@@
Figure 2: Supersymmetric RG flows from $N=2$ SCFT in five dimensions to $N=2$ SCFT in three dimensions from $ISO(3)\times U(1)$ $F(4)$ gauged supergravity.

### v3

@@ Abstract
We study holographic solutions describing RG flows across dimensions from five-dimensional $N=2$ SCFT to SCFTs in three and two dimensions using matter-coupled $F(4)$ gauged supergravity with $ISO(3)\times U(1)$ gauge group. By performing topological twists implemented by $SO(2)$, $SO(2)\times SO(2)$ and $SO(3)$ gauge fields, we find a number of $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ solutions preserving eight and four supercharges. These geometries are identified with conformal fixed points in three and two dimensions, respectively. The corresponding flow solutions are obtained numerically and can be interpreted as black membranes and black strings in asymptotically $AdS_{6}$ space. In addition, we have constructed charged domain wall solutions with $AdS_{2}\times S^{3}$ and $AdS_{3}\times S^{2}$ slicing. These solutions are supported by a non-vanishing two-form field and describe conformal line and surface defects within the $N=2$ SCFT in five dimensions, respectively. All of these solutions can be uplifted to type IIB theory via consistent truncations on $S^{2}\times\Sigma$ with $\Sigma$ being a Riemann surface.
@@ 1 Introduction
Five-dimensional field theories possess interesting non-trivial conformal fixed points at infinite coupling constants [1, 2, 3]. These fixed points are interesting in their own right due to the fact that they realize the only possible $N=2$ superconformal symmetry in five dimensions [4]. The essentially strongly-coupled nature and the lack of a Lagrangian formulation make understanding these field theories and the corresponding fixed points a non-trivial task. Up to now, it is well-known that the AdS/CFT correspondence [5, 6, 7] offers a new framework to gain more insights into the dynamics of strongly-coupled field theories via gravity duals. In AdS6/CFT5 correspondence, studies along this line have led to many interesting results, see for example [8]-[43]. These results, giving rise to various insights into different aspects of the five-dimensional superconformal field theories (SCFTs), have been obtained from both ten-dimensional string theories and effective six-dimensional gauged supergravities. Among these, holographic descriptions of twisted compactifications to lower-dimensional SCFTs [27, 26, 30, 31, 32, 33], as initiated in [44], and different types of conformal defects [34, 35, 36, 37, 38, 39, 40, 41, 42, 43] have recently attracted much attention.
In this paper, we are interested in solutions describing twisted compactifications of five-dimensional $N=2$ SCFT to three and two dimensions and conformal line and surface defects in the framework of six-dimensional $N=(1,1)$ gauged supergravity, usually called $F(4)$ gauged supergravity. All the previous works on these types of solutions have considered only pure $F(4)$ gauged supergravity [45] with $SU(2)$ gauge group or a simple extension to matter-coupled theory with $SU(2)\times SU(2)$ gauge group [46, 47]. The former can be embedded in massive type IIA theory using the result of [48] while the higher-dimensional origin of the latter is currently unknown. This renders many interesting holographic solutions, including a non-trivial $AdS_{6}$ vacuum with $SU(2)_{\textrm{diag}}$ symmetry identified in [24], not fully useful in the holographic context due to the lack of embedding in string/M-theory. We will partially fill this gap by considering $F(4)$ gauged supergravity coupled to four vector multiplets with $ISO(3)\times U(1)$ gauge group. This gauged supergravity and a truncation to three vector multiplets with $ISO(3)$ gauge group have been shown in [49] to arise from consistent truncations of type IIB theory on $S^{2}\times\Sigma$ for a Riemann surface $\Sigma$.
@@
For solutions describing conformal defects within $N=2$ SCFT in five dimensions, we will study solutions dual to line and surface defects preserving conformal symmetry in one and two dimensions. These solutions will be obtained by taking the metric ansatz in the form of $AdS_{2}\times S^{3}$- and $AdS_{3}\times S^{2}$-sliced domain walls. As we will see, the results found here and in [50] indicate that many aspects of the solutions are very similar to those previously found in compact $SU(2)$ and $SU(2)\times SU(2)$ gauge groups [43]. Although the structure of solutions is not rich as in the case of compact $SU(2)\times SU(2)$ gauge group due to the lack of a non-trivial $AdS_{6}$ vacuum in addition to the trivial one at the origin of the scalar manifold, many properties and even the form of the solutions are very similar. This is in agreement with the general expectation that solutions of gauged supergravity that has no known higher-dimensional origins might still capture some insights into the general aspect of strongly-coupled dynamics of the dual field theories.
The paper is organized as follows. In section 2, we review the general structure of matter-coupled $F(4)$ gauged supergravity as constructed in [46, 47]. We then consider the case of four vector multiplets and $ISO(3)\times U(1)$ gauge group and present the supersymmetric $AdS_{6}$ vacuum studied in [50]. In section 3, we study supersymmetric $AdS_{4}\times\Sigma$ and $AdS_{3}\times\mathcal{M}_{3}$ solutions using the domain wall ansatz with $\mathbb{R}^{1,2}\times\Sigma$ and $\mathbb{R}^{1,1}\times\mathcal{M}_{3}$ slicing and implementing topological twists by turning on $SO(2)\times U(1)$ and $SO(3)$ gauge fields. A number of $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ backgrounds including numerical solutions interpolating between these geometries and the supersymmetric $AdS_{6}$ vacuum will also be given. We then move to the study of solutions describing conformal line and surface defects in section 4. In this case, the metric ansatz takes the forms of $AdS_{2}\times S^{3}$- and $AdS_{3}\times S^{2}$-sliced domain walls, respectively. Unlike the case of twisted compactifications in which the two-form field from the supergravity multiplet can be consistently truncated out, we consider solutions with a non-vanishing two-form field and set all the gauge fields to zero. We find solutions describing line and surface defects due to a source of a dimension-$4$ operator dual to the two-form field. Finally, we give some conclusions and comments in section 5.
@@ 2 Matter-coupled $F(4)$ gauged supergravity with $ISO(3)\times U(1)$ gauge group
$e^{-1}\mathcal{L}$ $=$ $\frac{1}{4}R-\partial_{\mu}\sigma\partial^{\mu}\sigma-\frac{1}{4}P^{I\alpha}_{\mu}P^{\mu}_{I\alpha}-\frac{1}{8}e^{-2\sigma}\mathcal{N}_{\Lambda\Sigma}\widehat{F}^{\Lambda}_{\mu\nu}\widehat{F}^{\Sigma\mu\nu}-\frac{3}{64}e^{4\sigma}H_{\mu\nu\rho}H^{\mu\nu\rho}$
@@
$\widehat{F}^{\Lambda}=F^{\Lambda}-m\delta^{\Lambda 0}B,\qquad F^{\Lambda}=dA^{\Lambda}-\frac{1}{2}f^{\phantom{\Sigma}\Lambda}_{\Sigma\phantom{\Lambda}\Gamma}A^{\Sigma}\wedge A^{\Gamma},\qquad H=dB,$ (5)
where $f^{\phantom{\Sigma}\Lambda}_{\Sigma\phantom{\Lambda}\Gamma}$ are structure constants of the gauge group. It is also useful to note the convention on differential forms used in [46, 47]
@@
$\textrm{and}\qquad H_{\mu\nu\rho}$ $=$ $\frac{1}{3}(\partial_{\mu}B_{\nu\rho}+\partial_{\nu}B_{\rho\mu}+\partial_{\rho}B_{\mu\nu})$ (7)
@@
with ${\sigma^{rA}}_{B}$ being Pauli matrices. $SU(2)_{R}$ fundamental indices $A,B$ are raised and lowered by $\epsilon^{AB}$ and $\epsilon_{AB}$ with the convention $T^{A}=\epsilon^{AB}T_{B}$ and $T_{A}=T^{B}\epsilon_{BA}$. In this paper, we will use the values of $\epsilon^{12}=\epsilon_{12}=-1$. The covariant derivative of $\epsilon_{A}$ is defined by
@@
$Q_{\mu AB}=\frac{i}{2}\sigma^{r}_{AB}\left[\frac{1}{2}\epsilon^{rst}\Omega_{\mu st}-i\gamma_{7}\Omega_{\mu r0}\right].$ (19)
In this paper, we are interested in the case of $n=4$ vector multiplets leading to +×$\mathbb{R}^{+}\times SO(4,4)$ global symmetry. The gauge group of interest here is $ISO(3)\times U(1)$ with $ISO(3)$ embedded in $SO(4,3)\subset SO(4,4)$, and the $U(1)$ factor is the abelian gauge symmetry associated with the fourth vector multiplet as in the ungauged theory. The structure constants corresponding to the $ISO(3)\sim SO(3)\ltimes\mathbf{R}^{3}$ factor are given by
@@
with indices $I,J=1,2,3,4$ split as $I=(\bar{r},4)$ for $\bar{r},\bar{s}=1,2,3$. The resulting gauge generators are given by $X_{\Lambda}=(X_{0},X_{r},X_{\bar{r}})$ with $X_{0}=0$ while $X_{r}$ and $X_{r}-X_{\bar{r}}$ respectively generate the $SO(3)$ compact subgroup and three abelian translations $\mathbf{R}^{3}$ transforming as adjoint representation of $SO(3)$, see more detail in [50].
This gauged supergravity admits a supersymmetric $AdS_{6}$ vacuum at the origin of the $SO(4,4)/SO(4)\times SO(4)$ scalar manifold. This vacuum is given by, for more detail see [50],
@@
$V_{0}$ and $L$ denote the cosmological constant and the $AdS_{6}$ radius, respectively. As usual, we can choose $g=3m$ by shifting the value of the dilaton to $\sigma=0$ at the vacuum. Furthermore, this choice is also required by the embedding of six-dimensional gauged supergravity to type IIB theory given in [49]. With the value of the coupling constant $g=3m$, we find
@@
in which we have chosen $m>0$. This $AdS_{6}$ vacuum is dual to an $N=2$ SCFT in five dimensions.
@@ 3 Twisted compactifications of $N=2$ SCFT from ISO(3)$\times$×U(1) $F(4)$ gauged supergravity
In this section, we consider supersymmetric $AdS_{4}\times\Sigma$ and $AdS_{3}\times\mathcal{M}_{3}$ solutions with $\Sigma$ and $\mathcal{M}_{3}$ being a Riemann surface and a constant curvature $3$-manifold, respectively. These solutions are dual to conformal fixed points in three and two dimensions of the five-dimensional $N=2$ SCFT dual to the supersymmetric $AdS_{6}$ vacuum mentioned in the previous section. The solutions interpolating between these geometries and the $AdS_{6}$ vacuum are then interpreted as RG flows across dimensions from five dimensions to these fixed points in lower dimensions. In order to preserve some amount of supersymmetry, we need to perform a topological twist along $\Sigma$ and $\mathcal{M}_{3}$ by turning on suitable gauge fields to cancel spin connections on these manifolds. The lower-dimensional fixed points then arise from twisted compactifications of the $N=2$ SCFT in five dimensions on $\Sigma$ and $\mathcal{M}_{3}$. Moreover, the solutions can also be interpreted as black membranes and black strings in asymptotically $AdS_{6}$ space with near horizon geometries given by $AdS_{4}\times\Sigma$ and $AdS_{3}\times\mathcal{M}_{3}$.
@@ 3.1 Compactifications to three dimensions
$ds^{2}=e^{2f(r)}dx_{1,2}^{2}+dr^{2}+e^{2h(r)}(d\theta^{2}+F_{\kappa}(\theta)^{2}d\phi^{2})$ (23)
with $dx^{2}_{1,2}$ being a flat Minkowski metric in three dimensions with the signature (++)(-++)( - + + ). The function $F_{\kappa}(\theta)$ is defined by
@@
Denoting the six-dimensional coordinates by $x^{\mu}=(x^{a},r,\theta,\phi)$, for $a=0,1,2$, we can choose the following choice of vielbein
@@
The constants $a$ and $b$ correspond to the magnetic charges on $\Sigma$. With this ansatz for the gauge fields, it is consistent to set the two-form field to zero. A straightforward computation of the composite connection gives
$Q_{\mu AB}=-\frac{i}{2}gA_{\mu}^{3}\sigma^{3}_{AB}\,.$ (31)
@@
$0=\frac{1}{2}\frac{F^{\prime}_{\kappa}(\theta)}{F_{\kappa}(\theta)}e^{-h}\gamma_{\hat{\phi}\hat{\theta}}\epsilon_{A}-\frac{i}{2}ga\frac{F^{\prime}_{\kappa}(\theta)}{F_{\kappa}(\theta)}e^{-h}\sigma^{3}_{AB}\epsilon^{B}+\ldots\,.$ (32)
The remaining terms involving $h^{\prime}$, scalars and gauge fields take a similar form as in $\delta\psi_{A\hat{\theta}}$ variation. These will eventually lead to a BPS equation for $h(r)$ after the cancellation of the two terms given in (32).
We now perform a topological twist by imposing a projector
@@
With all these results, we are ready to analyze the BPS conditions arising from setting supersymmetry transformations of fermions to zero. With only the $(\theta\phi)$-component of the field strength tensors non-vanishing, the field equations of gauge fields require $\phi_{3}=0$. Furthermore, consistency of the BPS equations also require $\phi_{0}=\phi_{4}=0$. By imposing another projector of the form
@@
It can be readily verified that these equations are compatible with the corresponding field equations. The sign choice in (36) has been chosen such that the $AdS_{6}$ vacuum appears at $r\rightarrow\infty$. In addition, for $a=b=0$ which effectively sets all the gauge fields to zero, these equations reduce to those considered in [50] for finding holographic RG flow solutions.
@@
It should be pointed out that, from equation (39), $AdS_{4}$ fixed points with $\phi_{5}=0$ are possible only for $b=0$ which implies $A^{7}=0$. Another possibility to satisfy equation (39) for $\phi_{5}=0$ is given by setting $\kappa=0$, but this does not lead to any $AdS_{4}$ fixed points. From equation (43) and $g=3m$ with $m>0$, it is clearly seen that only $\kappa=-1$ is possible leading to an $AdS_{4}\times H^{2}$ solution. This solution can also be interpreted as an $AdS_{4}\times H^{2}$ solution of $F(4)$ gauged supergravity coupled to three vector multiplets with $ISO(3)$ gauge group since all the fields from the fourth vector multiplet have been truncated out.
Another $AdS_{4}\times\Sigma$ solution is given by
@@
We have also chosen $a>b$ in these equations. As in the previous solution, this fixed point exists only for $\kappa=-1$ leading to another $AdS_{4}\times H^{2}$ solution. These solutions should be dual to conformal fixed points in three dimensions arising from twisted compactifications of $N=2$ SCFT in five dimensions, dual to the $AdS_{6}$ vacuum, on a two-dimensional hyperbolic space $H^{2}$. The solutions interpolating between these backgrounds then describe holographic RG flows from five to three dimensions. On the other hand, these solutions can also be regarded as black membranes with the near horizon geometry given by $AdS_{4}\times H^{2}$ in asymptotically $AdS_{6}$ space.
Before giving interpolating solutions between $AdS_{6}$ and $AdS_{4}\times H^{2}$ geometries, we note the unbroken supersymmetry preserved by the aforementioned solutions. At the $AdS_{4}$ fixed points with constant scalars, the $\gamma_{r}$ projector given in (36) is irrelevant. Accordingly, the $AdS_{4}\times H^{2}$ solutions are half-BPS, due to the projector (33), and preserve eight supercharges corresponding to $N=2$ superconformal symmetry in three dimensions. On the other hand, the flow solutions interpolating between $AdS_{6}$ and $AdS_{4}\times H^{2}$ vacua preserve only four supercharges due to the two projectors given in (33) and (36).
@@
Figure 1: Supersymmetric RG flows from $N=2$ SCFT in five dimensions (dual to the $AdS_{6}$ vacuum (21)) to $N=2$ SCFT in three dimensions (dual to the $AdS_{4}\times H^{2}$ vacuum (43)) from $ISO(3)$ $F(4)$ gauged supergravity with $g=3m$ and $m=\frac{1}{4}$ (red), $m=\frac{1}{2}$ (green), $m=1$ (blue).
@@
Figure 2: Supersymmetric RG flows from $N=2$ SCFT in five dimensions (dual to the $AdS_{6}$ vacuum (21)) to $N=2$ SCFT in three dimensions (dual to the $AdS_{4}\times H^{2}$ vacuum (44)) from $ISO(3)\times U(1)$ $F(4)$ gauged supergravity with $g=3m$, $m=\frac{1}{4}$ and $b=-1$ (red), $b=1$ (green), $b=2$ (blue).